Nathan Detroit owns 100 shares of the drink company Monster Beverage​ Corp., which he purchased for ​$120 per share. Nathan read in the Wall Street Journal that the​ company's board of directors had voted to split the stock 5​-for-1. Just before the stock​ split, Monster Beverage shares were trading for ​$129.02. Answer the following questions about the impact of the stock split on his holdings and taxes. Nathan is in the 22 % federal income tax bracket.

a.  How many shares of Monster Beverage will Nathan own after the stock​ split?

b.  Immediately after the​ split, what do you expect the value of Monster Beverage to​ be?

c.  Compare the total value of​ Nathan's stock holdings before and after the split. What do you​ find?

d.  Does Nathan experience a gain or loss on the stock as a result of the 5​-for-1 ​split?

e.  What is​ Nathan's tax liability from the​ event?
Answer #1

a.

Number of shares held by Nathan before split = 100

Number of Shares after split = 5(100) = 500 shares

b.

Value of Share after Split = 129.02/5 = $25.804

c.

Total Value of Nathan's stock holding will remain same before and after split.

100(129.02) = 500(25.804)

d.

Nathan doesn't experience any gain or loss due to split